OpenStudy (anonymous):

Medal During a ultrasound, the technician reveals that she is carrying quadruplets - 4 fraternal twins! Because there are four babies in the womb, the technician is unable to tell the gender of the babies. (a) Find the probability that the quadruplets will all be boys. Express your answer as a fraction or an exact decimal - do not round. (b) What is the probability that the couple will have 1 boy and 3 girls? (c) What is the probability that the oldest three children will be female?

OpenStudy (perl):

would you like help ?

OpenStudy (perl):

a) find probability of BBBB

OpenStudy (anonymous):

how

OpenStudy (anonymous):

yes i would like your help please

OpenStudy (perl):

the probability of getting a boy is 1/2

OpenStudy (anonymous):

ok is that part a?

OpenStudy (perl):

that is part of part a) :)

OpenStudy (perl):

it is asking for BBBB

OpenStudy (anonymous):

ya i understand

OpenStudy (perl):

ok,

OpenStudy (perl):

for part b) we want BGGG (and there are others

OpenStudy (anonymous):

3/4

OpenStudy (perl):

b) BGGG GBGG GGBG GGGB

OpenStudy (perl):

wait, what did you get for part a) ?

OpenStudy (anonymous):

1/2

OpenStudy (perl):

thats incorrect

OpenStudy (perl):

P( BBBB ) = P(B ) * P(B) * P(B) * P(B)

OpenStudy (anonymous):

3/4

OpenStudy (perl):

1/2 * 1/2 * 1/2* 1/2

OpenStudy (anonymous):

4/8

OpenStudy (perl):

nope :)

OpenStudy (perl):

1/2 * 1/2 = 1/4 correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1/8

OpenStudy (perl):

ok, almost there

OpenStudy (anonymous):

ohhhh ya i c it yaa

OpenStudy (perl):

so you can use a similar idea (multiply probabilities) for part b)